PROCEDURES AND SUBROUTINES FOR ANALYSIS OF SPECTRAL DATA, OF TYPES BOTH FRE-QUENCY AND INTENSITY, OF DIATOMIC MOLECULAR SPECIES

\underline{J. F. OGILVIE}, Centre for Experimental and Constructive Mathematics, Department of Mathematics, Simon Fraser University, 8888 University Drive, Burnaby, BC V5A 1S6 Canada.